RC meeting and Run Plan, Tuesday November 1st

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room

Useful Links

Current Status

Polarized Target

  • Current Polarization: N/A
  • Requesting hourly screenshots of Target Strip Charts (and keep an eye on these during shift) - Should be removed after configuration change ended

Run Status

  • Hall was in controlled access last night and this morning.
    • Experts performing a series of measurements this afternoon and tonight/overnight.
    • More tests of the new VME crate (BPMs and more)
    • Jessie and others did work to prepare for major reconfiguration tomorrow
    • Other minor work: over interlock box replaced, pulse generator for GRINCH
  • Other Subsystem updates?

Future Plans

  • Hall in power permit for magnet-on target measurements. Will go back to CA later this afternoon: current estimate is around 4pm
    • Plan on more measurements in power permit overnight, as needed
  • Tomorrow morning, move to Restricted Access
    • Radcon will come in ~7:30 to bag up contaminated items in target enclosure; prepare for work
    • Target group will prepare target enclosure for field measurements and then make the measurements
    • During some of this time, the BB/SBS magnets will be ON during the Controlled Access (under special test procedures). Check with RC before going down
  • Still to do/schedule:
    • Remaining BPM crate replacement work? - Keith
  • Any other access needed tonight, tomorrow, or Wednesday?
  • MCC will set up separator on Wednesday (for 3-pass delivery)
    • Most likely scenario is that we're ready to request beam on Friday; details depend on timing of work in the hall.

Other Topics

Experimental progress (for MCC planning meeting): Based on Run Coordinator instructions (new this week)

  • 1.1% of accumulated Figure of Merit (charge, corrected for target polarization squared) is 5.19 C out of 476 C desired:
  • 27% Experimental run time (38 out of 141 days)
  • Notes:
    • Raw charge is 13.55 C [2.8% of desired uncorrected charge]
    • First 15 days had no beam [40% of our experiment time so far]
    • Desired charge neglects all time for commissioning, configuration changes, etc... --> expect it to grow much faster as run gets going
  • Is this what we want to present??

Current Run Plan

Configuration change.

Need updates for the 'standard' run plan below ------

NMR measurement Restore beam

Run Plan 1.) He-3 runs, at 45uA of beam current. Use ps3==0, ps5=12, all others -1. Take four runs that should each last for one hour. 2.) Move the target to the pickup coil position, and the TO should run an NMR sweep to measure the polarization. 3.) Return to step 1.)

  • Keep an eye out for drifting BPM positions
  • Log all DAQ crashes in the excel worksheet along with the reason for the crash
  • If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: [1], If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
  • Cross check the rate with what is expected: [2]
  • Cross check with the expected ion chamber readings: [3]

Notices in Effect for Shift Crew

Current Issues

  • Shift Openings: Monday (morning) OWL, Tuesday OWL and SWING. [4]

Planned Accesses to the Hall

Special Requests